OpenStudy (anonymous):

anybody good at optimization? i am stuck on an online answer problem and the answer key has a mistake in it

OpenStudy (anonymous):

post it, you will get lots of answers i am sure

OpenStudy (anonymous):

A window is being built and the bottom is a rectangle and the top is a semicircle. If there is 12 meters of framing materials what must the dimensions of the window be to let in the most light?

OpenStudy (anonymous):

to let in the most lite it should have the maximum area A=lb we have perimeter= 2b+l+pi*l/2=12 substitute b from above equation in the area equation and maximise area with respect to l you get length from that find b

OpenStudy (anonymous):

b=6-(2+pi)*l/4 A=6l-(2+pi)l^2/4

OpenStudy (anonymous):

thank you!